OpenStudy (anonymous):

write using exponents (-4)(-4)= please help

OpenStudy (anonymous):

\[\left( -4 \right)^{2}\]

OpenStudy (anonymous):

you need the parentheses around a negative number

OpenStudy (anonymous):

is the answer (-4) with the little 2

OpenStudy (anonymous):

yes which equals 16

OpenStudy (anonymous):

this is confuseing

OpenStudy (anonymous):

ok so you multiply the 2 fours/

OpenStudy (anonymous):

yes (-4)^2 = (-4)(-4) which was the original problem. =16

OpenStudy (anonymous):

ok so when i write the problem using exponents i write it like (-4) ^2= (-4)(-4) that si what i would write?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

ok i am sorry i really am not good in math i really appreciate your help.

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

i have another one i am haveing a problem with could you help me?

OpenStudy (anonymous):